OpenStudy (anonymous):

Calculate δf/δx and δf/δy for the function: f^2 +cos(xy)=f How do I do partial differentiation in this form? (Normally it is f(x,y)=...)

OpenStudy (anonymous):

woah, thats funky.

OpenStudy (loser66):

to me, f^2 -f + cos(xy) =0 2f df/dx -df/dx -y sin(xy)=0 df/dx(2f-1) =y sin(xy) df/dx = \(\dfrac{y sin(xy)}{2f-1}\)

OpenStudy (anonymous):

I see, so, df/dy= \[\frac{ xsinxy }{ 2f-1 }\]\

OpenStudy (loser66):

:)

OpenStudy (anonymous):

Thanks for your help in the previous question btw. Although it was not the actual answer, but I figured it out.